OpenStudy (anonymous):

Solve the inequality: y - 2 < 5 or y - 2 > 5 Then describe the graph of the solution. (Will the starting points be open or closed? Where are the starting points placed on the number line? Which way will each arrow point?

OpenStudy (anonymous):

y= 7 for y-2<5 put y= 0 -2< 5 y= 6.9 4.9<5 true put y= 7.1 5.1<5, false

OpenStudy (anonymous):



OpenStudy (anonymous):

But how on a number line would this look? The starting point would be closed, right? but where would the point start and end at?

OpenStudy (anonymous):

its y < 7 start at 6.99999999999 ends at - infinity